OpenStudy (anonymous):

6. Solve the equation. Show your work. Check your answer. 3x – 7 = 5x + 19 can someone help me with this

OpenStudy (misssunshinexxoxo):

Simplifying 3x + -7 = 5x + 19 Reorder the terms: -7 + 3x = 5x + 19 Reorder the terms: -7 + 3x = 19 + 5x Solving -7 + 3x = 19 + 5x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-5x' to each side of the equation. -7 + 3x + -5x = 19 + 5x + -5x Combine like terms: 3x + -5x = -2x -7 + -2x = 19 + 5x + -5x Combine like terms: 5x + -5x = 0 -7 + -2x = 19 + 0 -7 + -2x = 19 Add '7' to each side of the equation. -7 + 7 + -2x = 19 + 7 Combine like terms: -7 + 7 = 0 0 + -2x = 19 + 7 -2x = 19 + 7 Combine like terms: 19 + 7 = 26 -2x = 26 Divide each side by '-2'. x = -13 Simplifying x = -13

OpenStudy (kash_thesmartguy):

Simplifying 3x + -7 = 5x + 19 Reorder the terms: -7 + 3x = 5x + 19 Reorder the terms: -7 + 3x = 19 + 5x Solving -7 + 3x = 19 + 5x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-5x' to each side of the equation. -7 + 3x + -5x = 19 + 5x + -5x Combine like terms: 3x + -5x = -2x -7 + -2x = 19 + 5x + -5x Combine like terms: 5x + -5x = 0 -7 + -2x = 19 + 0 -7 + -2x = 19 Add '7' to each side of the equation. -7 + 7 + -2x = 19 + 7 Combine like terms: -7 + 7 = 0 0 + -2x = 19 + 7 -2x = 19 + 7 Combine like terms: 19 + 7 = 26 -2x = 26 Divide each side by '-2'. x = -13 Simplifying x = -13

OpenStudy (anonymous):

Whoa your smart XD Thanks

OpenStudy (misssunshinexxoxo):

@Kash_TheSmartGuy No

OpenStudy (kash_thesmartguy):

Oh, I thought so, sorry :(

OpenStudy (misssunshinexxoxo):

No worries, pleasure to meet you on OS

OpenStudy (anonymous):

Thanks Kash

OpenStudy (anonymous):

You're most welcome :) going to try to catch up to what you did there.